# and Machine Learning

### v3

@@ 1 Introduction
Like all things of fundamental importance, image denoising is easy to describe, and very difficult to do well in practice. It is therefore not surprising that the field has been around since the beginning of the modern scientific and technological age - for as along as there have been sensors to record data, there has been noise to contend with.
@@
$\mathbf{\hat{x}(\alpha)}=f(\mathbf{x};\alpha)\approx\mathbf{u}.$ (2)
where $\alpha(\sigma^{2})$ is a monotonic function of the noise variance, and therefore controls the ”strength” of the denoiser.
As the description above indicates, a denoiser is not a single operator but a family of bounded [^2] We assume all images are in the numerical range [$0,1$]. In practice, an $8$-bit image would have values in [$0,255$] range. maps $f(\mathbf{x},\alpha):\left[0,1\right]^{N}\rightarrow\left[0,1\right]^{N}$. We expect ”good” denoisers to have certain naturally desirable properties, which alas in practice, many do not. For the sake of completeness, and as a later guide for how to design good denoisers, we call a denoiser ideal if it satisfies the following properties:
@@
$f(\mathbf{x},0)=\mathbf{x},\;\;\;\;\;\;\forall\mathbf{x}$ (3)
@@
(Conservation) An ideal denoiser has a symmetric Jacobian
$\nabla f(\mathbf{x},\alpha)=\nabla f(\mathbf{x},\alpha)^{T}$ (4)
@@
$f(\mathbf{x},\alpha)=\nabla\mathcal{E}(\mathbf{x},\alpha)$ (5)
for some scalar-valued, differentiable (”potential” or ”energy”) function $\mathcal{E}(\mathbf{x},\alpha)$. This also means that the ideal denoiser defines a conservative vector field.
To convey some intuition for this property, consider the linear case. When a denoiser is linear: $f(\mathbf{x},\alpha)=W(\alpha)\mathbf{x}$, we always require the matrix $W(\alpha)$ to be row stochastic (meaning the rows sum to $1$) in order to preserve the mean local brightness. Ideally, we also require $W(\alpha)$ to be symmetric [1], which has the added advantage that the denoiser is admissible [2] in the mean-square sense. Property $2$ extends these notions to more general nonlinear denoisers  [^3] It’s worth noting that the combination of symmetric and row-stochastic implies that $W(\alpha)$ is doubly-stochastic..
@@
The conservation Property 2 guarantees that the ideal denoiser is the gradient of a scalar field. When the Hessian of the scalar field is bounded, this also implies that $f(\mathbf{x},\alpha)$ is a Lipschitz map with some constant $M(\alpha)$:
$\|f(\mathbf{x},\alpha)-f(\mathbf{y},\alpha)\|\leq M(\alpha)\|\mathbf{x}-\mathbf{y}\|$ (6)
@@
$g(\mathbf{x},\underline{\alpha})=\sum_{k=0}^{N}a_{k}\>f(\mathbf{x},\alpha_{k})\;\;\;\;\;\;\;\;\;\text{with}\;\;\;\;\;\sum_{k=0}^{N}\alpha_{k}=1$ (7)
it is easy to verify that Properties $1$ and $2$ are satisfied.
  Summary: Ideal denoisers satisfy:
@@
A note on this work: Rather than a survey of image denoising, this work focuses on defining ideal denoisers, their properties, and their connections to statistical theory and machine learning. We then demonstrate how these powerful components can serve as building blocks in various applications. Readers interested in a historical overview of image denoising are encouraged to consult the excellent resources in  [6, 7, 8, 9, 10].
@@ 2 Denoising as a Natural Decomposition
One of the remarkable aspects of well-behaved (even if not ideal) denoising operators is that we can employ them to easily produce a natural multi-scale decomposition of an image, with perfect reconstruction property [^5] This is similar in spirit to a multiscale decomposition, except that there is no decimation, and the filters are nonlinear here.. To start, consider a denoiser $f(\mathbf{x},\alpha)$. We can write the obvious relation:
$\mathbf{x}=f(\mathbf{x},\alpha)+\left[\mathbf{x}-f(\mathbf{x},\alpha)\right]$ (8)
The first term on the right-hand side is a smoothed (or denoised) version of $\mathbf{x}$, whereas the second term in the brackets is the residual $r_{0}(\mathbf{x},\alpha)=\mathbf{x}-f(\mathbf{x},\alpha)$ which is an ostensibly ”high-pass” version. Next, we can apply the same decomposition repeatedly to the already-denoised components [^6] To simplify the exposition, we use the same denoiser and the $\alpha$ at each step, but this is not necessary.:
@@
$=$ $f(f(\mathbf{x},\alpha),\alpha)+r_{1}(\mathbf{x},\alpha)+r_{0}(\mathbf{x},\alpha)$ (10)
$\vdots$ (11)
$=$ $f^{n}(\mathbf{x},\alpha)+\sum_{k=0}^{n-1}r_{k}(\mathbf{x},\alpha)$ (12)
@@
$g_{a}(\mathbf{x},\alpha)=\beta_{n}f^{n}(\mathbf{x},\alpha)+\sum_{k=0}^{n-1}\beta_{k}r_{k}(\mathbf{x},\alpha).$ (13)
This approach was generalized and used in a practical setting in [14, 15] to produce a wide variety of image processing effects, built on a base of well-established (at the time) non-local means denoisers. This is illustrated in Figure 1. More recently, in [16] the authors used a similar decomposition to create a zero-shot method to control each individual component of the decomposition through diffusion model sampling.
@@
The concept of breaking down an image into layers of varying detail is closely related to the architecture of Residual Neural Networks [17] (ResNets). Both share the principle that it’s simpler to model/learn residual mappings (the difference between the input and desired output) than to model/learn the complete transformation directly. While traditional deep neural networks try to learn this complex mapping in one go, ResNets use “skip connections” that allow the network to bypass layers, adding the original input to a later layer’s output. Let $H(\mathbf{x})$ be the desired complex mapping and $\mathbf{x}$ is the input, a ResNet layer attempts to learn a residual function $F(\mathbf{x})$ such that:
@@
The skip connection ensures that the original input $\mathbf{x}$ is preserved and added back to the output. Note the connection to (8), where the residual term is exactly $-F(\mathbf{x})$. This decomposition and the use of skip connections simplify the network’s task, making optimization easier and mitigating the vanishing gradient problem that can hinder deep network training. Additionally, the preservation of the original input or its smooth approximation through skip connections ensures important information isn’t lost as data travels through the network [^7] There are actually ways of ensuring invertibility of ResNets, see e.g.,[18]. ResNets have been a major breakthrough in deep learning, enabling the training of much deeper networks and achieving state-of-the-art performance on image recognition tasks, with the concept of residual learning now being applied to other domains beyond image processing.
@@ 3.1 Bayesian Denoisers
where $\phi(\mathbf{u})=-\log P(\mathbf{u})$ is the negative log-prior on the space of ”clean” images, and $\alpha$ is proportional to the noise variance. It would appear that the MAP denoiser does not have a closed form. However, the expression (16) is also known in the optimization literature as a proximal operator [21] when $\phi$ is convex, quasi-convex, or a difference of convex functions. It is well-known [22] that to every proximal operator $f$ there corresponds a convex scalar-valued function $\psi$ such that $f=\nabla\psi$.
@@
where $\phi_{\alpha}$ is a smoothed version of $\phi$ called its Moreau envelope [21, 23]. As we will see below, the MMSE estimate shares a very similar form.
@@
In contrast to MAP denoising, MMSE denoising relies on the mean of the posterior density,
@@
While the MMSE expectation integral is generally difficult or impossible to evaluate directly, a key result known as Tweedie’s formula [24, 25, 26] enables us to write the expression for MMSE also in the form of the gradient of a scalar function:
@@
Data-driven MMSE denoisers: The typical modern supervised approach to image denoising is to train a deep neural network with pairs of clean and noisy images, where the noise is often modeled as additive white and Gaussian(AWGN). Let’s assume we have image pairs $(\mathbf{u},\mathbf{x})\sim P(\mathbf{u},\mathbf{x})$ where $\mathbf{u}$ represents a clean image, and $\mathbf{x}$ is the noisy observation obtained by adding AWGN with a known standard deviation to $\mathbf{u}$: $\mathbf{x}=\mathbf{u}+\mathbf{e}$, where $\mathbf{e}\sim\mathcal{N}(0,\sigma^{2})$.
A typical regression approach would predict $\mathbf{u}$ directly from $\mathbf{x}$ using a trained model $\hat{\mathbf{u}}(\mathbf{x})=F_{\theta}(\mathbf{x})\approx\mathbf{u}$, by minimizing the expected reconstruction error:
$\min_{\theta}\mathbb{E}_{\mathbf{u},\mathbf{e}}\|F_{\theta}(\mathbf{u}+\mathbf{e})-\mathbf{u}\|^{p}_{p}\approx\min_{\theta}\sum_{i}\|F_{\theta}(\mathbf{u}^{i}+\mathbf{e}^{i})-\mathbf{u}^{i}\|^{p}_{p}.$
@@
As mentioned earlier, the MMSE denoiser is the average of all plausible clean signals given the noisy observation. This averaging can lead to a loss of details and a blurry appearance, especially when the noise level is high.
@@ 3.2 Energy-based Denoisers
We’ve seen that both MMSE and MAP estimators are of the form $f(\mathbf{x},\alpha)=\mathbf{x}-\alpha\nabla\phi_{\alpha}(\mathbf{x})$ where $\phi_{\alpha}(\mathbf{x})$ is some smoothed version of $\phi$ – differently smoothed in each case. These denoisers are also special cases of a general ”energy-based” formulation [28]:
$f(\mathbf{x},\alpha)=\nabla\mathcal{E}(\mathbf{x},\alpha)$ (21)
@@
$\mathcal{E}(\mathbf{x},\alpha)=\frac{1}{2}\|\mathbf{x}\|^{2}-\alpha\phi_{\alpha}(\mathbf{x})$ (22)
@@
$\nabla f(\mathbf{x},\alpha)=\mathcal{H}\left[\mathcal{E}(\mathbf{x},\alpha)\right]$ (23)
@@
3.2.1 Approximation of Energy-based Denoisers
The energy-based formulation of denoisers provides a natural mechanism for principled empirical design of denoisers. This approach turns out to be consistent with the well-established empirical Bayes approach as well.
@@ 3.2.1 Approximation of Energy-based Denoisers
$s(\mathbf{x},\alpha)=\nabla_{\mathbf{x}}\log P(\mathbf{x},\alpha)=-\nabla_{\mathbf{x}}\mathcal{E}(\mathbf{x},\alpha).$ (25)
@@
$s(\mathbf{x},\alpha)=-\nabla_{\mathbf{x}}\mathcal{E}(\mathbf{x},\alpha)=-\nabla\mathcal{E}(0,\alpha)-\mathbf{L}(\mathbf{a},\alpha)\mathbf{x}.$ (27)
Tweedie’s formula implies that the MMSE denoiser has the form:
$f(\mathbf{x},\alpha)$ $=\mathbf{x}-\alpha\>s(\mathbf{x},\alpha)$ (28)
$=\mathbf{x}+\alpha\nabla\mathcal{E}(0,\alpha)+\alpha\mathbf{L}(\mathbf{a},\alpha)\mathbf{x}.$
@@
$f(\mathbf{x},\alpha)=\mathbf{W}(\mathbf{x},\alpha)\>\mathbf{x}$ (30)
@@
Though these facts were neither historically clarified, nor the original motivation for their development [29], denoisers of the form (30) have always been heuristic/empirical approximations to the MMSE. These denoisers were hugely popular and effective (e.g. [32, 11]) for decades before the more recent introduction of neural networks. More recent work by Scarvelis et al. [33] explores the use of a specific kernel approach to create a ”closed-form” diffusion model that operates directly on the training set, without the need for training a neural network.

### v4

@@ 1 Introduction
Like most things of fundamental importance, image denoising is easy to describe, and very difficult to do well in practice. It is therefore not surprising that the field has been around since the beginning of the modern scientific and technological age - for as along as there have been sensors to record data, there has been noise to contend with.
@@
$\mathbf{\hat{x}(\alpha)}=f(\mathbf{x};\alpha)\approx\mathbf{u},$ (2)
where $\alpha(\sigma^{2})$ is a monotonic function of the noise variance, and therefore controls the “strength” of the denoiser.
As the description above indicates, a denoiser is not a single operator but a family of bounded [^2] We assume all images are in the numerical range [$0,1$]. In practice, an $8$-bit image would have values in [$0,255$] range. maps $f(\mathbf{x},\alpha):\left[0,1\right]^{N}\rightarrow\left[0,1\right]^{N}$. We expect “good” denoisers to have certain naturally desirable properties, which alas in practice, many do not. For the sake of completeness, and as a later guide for how to design good denoisers, we call a denoiser ideal if it satisfies the following properties:
@@
$f(\mathbf{x},0)=\mathbf{x},\;\;\;\;\;\;\forall\mathbf{x}.$ (3)
@@
(Conservation) An ideal denoiser has a symmetric Jacobian [^3] Unless explicitly noted otherwise, $\nabla$ will mean $\nabla_{\mathbf{x}}$ throughout the paper.
$\nabla f(\mathbf{x},\alpha)=\nabla f(\mathbf{x},\alpha)^{T}.$ (4)
@@
$f(\mathbf{x},\alpha)=\nabla\mathcal{E}(\mathbf{x},\alpha),$ (5)
for some scalar-valued, differentiable (“potential” or “energy”) function $\mathcal{E}(\mathbf{x},\alpha)$. This also means that the ideal denoiser defines a conservative vector field [^4] We note that since the ideal denoiser can be expressed as the gradient of a scalar function, this leads directly to the path independence property of line integrals, which is the defining characteristic of a conservative vector fields..
To convey some intuition for this property, consider the linear case. When a denoiser is linear: $f(\mathbf{x},\alpha)=W(\alpha)\mathbf{x}$, we always require the matrix $W(\alpha)$ to be row-stochastic (meaning the rows sum to $1$) in order to preserve the mean local brightness. Ideally, we also require $W(\alpha)$ to be symmetric [1], which has the added advantage that the denoiser is admissible [2] in the mean-square sense. Property $2$ extends these notions to more general nonlinear denoisers  [^5] It’s worth noting that the combination of symmetric and row-stochastic implies that $W(\alpha)$ is doubly-stochastic..
@@
The conservation Property 2 guarantees that the ideal denoiser is the gradient of a scalar field.This also implies that $f(\mathbf{x},\alpha)$ is a Lipschitz map with some constant $M(\alpha)$:
$\|f(\mathbf{x},\alpha)-f(\mathbf{y},\alpha)\|\leq M(\alpha)\|\mathbf{x}-\mathbf{y}\|.$ (6)
@@
$g_{a}(\mathbf{x},\alpha)=\sum_{k=0}^{N}a_{k}\>f(\mathbf{x},\alpha_{k})\;\;\;\;\;\text{with}\;\;\;\;\;\sum_{k=0}^{N}a_{k}=1,$ (7)
it is easy to verify that Properties $1$ and $2$ are satisfied.   Summary: Ideal denoisers satisfy:
@@
A note on this work: Rather than a survey of image denoising, this work focuses on defining ideal denoisers, their properties, and their connections to statistical theory and machine learning. We then demonstrate how these powerful components can serve as building blocks in various applications. Readers interested in a historical overview of image denoising are encouraged to consult the excellent resources in  [6, 7, 8, 9, 10]. Our analysis specifically considers an additive white Gaussian noise model due to its broad applicability and relevance to the applications explored herein. A deeper examination of various noise models can be found in [11, 6, 12].
@@ 2 Denoising as a Natural Decomposition
One of the remarkable aspects of well-behaved (even if not ideal) denoising operators is that we can employ them to easily produce a natural multiscale decomposition of an image, with perfect reconstruction property [^7] This is similar in spirit to the classic multiscale decomposition in [13], except that there is no decimation, and the filters are nonlinear here.. To start, consider a denoiser $f(\mathbf{x},\alpha)$. We can write the obvious relation:
$\mathbf{x}=f(\mathbf{x},\alpha)+\left[\mathbf{x}-f(\mathbf{x},\alpha)\right].$ (8)
The first term on the right-hand side is a smoothed (or denoised) version of $\mathbf{x}$, whereas the second term in the brackets is the residual $r_{0}(\mathbf{x},\alpha)=\mathbf{x}-f(\mathbf{x},\alpha)$ which is an ostensibly “high-pass” version. Next, we can apply the same decomposition repeatedly to the already-denoised components [^8] To simplify the exposition, we use the same denoiser and the $\alpha$ at each step, but this is not necessary.:
@@
$=$ $f(f(\mathbf{x},\alpha),\alpha)+r_{1}(\mathbf{x},\alpha)+r_{0}(\mathbf{x},\alpha)$
$\vdots$
$=$ $f^{n}(\mathbf{x},\alpha)+\sum_{k=0}^{n-1}r_{k}(\mathbf{x},\alpha),$
@@
$g_{a}(\mathbf{x},\alpha,\beta)=\beta_{n}f^{n}(\mathbf{x},\alpha)+\sum_{k=0}^{n-1}\beta_{k}r_{k}(\mathbf{x},\alpha).$ (10)
This approach was generalized and used in a practical setting in [17, 18] to produce a wide variety of image processing effects, built on a base of well-established (at the time) non-local means denoisers. This is illustrated in Figure 1. More generally, given paired examples of input and desired output images ($\mathbf{x}_{i},\widehat{\mathbf{x}}_{i}$), one can construct a loss function such as shown below, where $d$ is a training loss, and $\mathcal{R}$ is a regularization term. By minimizing this loss, we can learn both the parameters $\alpha$ and $\beta$.
$\text{Loss}_{f}(\alpha,\beta)=\frac{1}{N}\sum_{i=1}^{N}d(\widehat{\mathbf{x}}_{i},g_{\beta}(\mathbf{x}_{i},\alpha))+\mathcal{R}(\alpha,\beta).$ (11)
Recently, in [19] the authors used a similar decomposition to create a zero-shot method to control each individual component of the decomposition through diffusion model sampling.
@@
The concept of breaking down an image into layers of varying detail is closely related to the architecture of Residual Neural Networks [20] (ResNets). Both share the principle that it’s simpler to model/learn residual mappings (the difference between the input and desired output) than to model/learn the complete transformation directly. While traditional deep neural networks try to learn this complex mapping in one go, ResNets use “skip connections” that allow the network to bypass layers, adding the original input to a later layer’s output. Letting $H(\mathbf{x})$ be the desired complex mapping and $\mathbf{x}$ the input, a ResNet layer attempts to learn a residual function $F(\mathbf{x})$ such that:
@@
The skip connection ensures that the original input $\mathbf{x}$ is preserved and added back to the output. Note the connection to (8), where the residual term is exactly $-F(\mathbf{x})$. This decomposition and the use of skip connections simplify the network’s task, making optimization easier and mitigating the vanishing gradient problem that can hinder deep network training [21]. Additionally, the preservation of the original input or its smooth approximation through skip connections ensures important information isn’t lost as data travels through the network [^9] There are actually ways of ensuring invertibility of ResNets, see e.g.,  [22].. ResNets have been a major breakthrough in deep learning, enabling the training of much deeper networks and achieving state-of-the-art performance on image recognition tasks, with the concept of residual learning now being applied to other domains beyond image processing.
@@ 3.1 Bayesian Denoisers
where $\phi(\mathbf{u})=-\log P(\mathbf{u})$ is the negative log-prior on the space of “clean” images, and $\alpha$ is proportional to the noise variance. It would appear that the MAP denoiser does not have a closed form. However, the expression (14) is also known in the optimization literature as a proximal operator [25, 26] when $\phi$ is convex, quasi-convex, or a difference of convex functions. It is well-known  [25, 27] that to every proximal operator $f$ there corresponds a convex scalar-valued function $\psi$ such that $f=\nabla\psi$.
@@
where $\phi_{\alpha}$ is a smoothed version of $\phi$ called its Moreau envelope [25, 28, 29, 30]. As we will see below, the MMSE estimate shares a very similar form.
@@
While Maximum A Posteriori (MAP) denoising seeks the most probable estimate of a clean signal given a noisy observation, MMSE denoising aims to find the estimate that minimizes the mean squared error (MSE) between the estimate and the true signal
$MSE(\hat{\mathbf{x}},\mathbf{u})=\mathbb{E}_{\mathbf{u},\mathbf{x}}\left[\|\hat{\mathbf{x}}-\mathbf{u}\|^{2}_{2}\right],$ (16)
where $\mathbf{u}$ is the true signal, $\mathbf{x}$ the noisy observation, and $\hat{\mathbf{x}}$ is the estimate of $\mathbf{u}$ given $\mathbf{x}$.
The Posterior Mean as the MMSE Estimator. A fundamental result in estimation theory is that the posterior mean, $\mathbb{E}\left[\mathbf{u}|\mathbf{x}\right]$, is the MMSE estimator. This can be shown by minimizing the MSE directly. Starting with the definition of MSE:
MSE $=\mathbb{E}_{\mathbf{u},\mathbf{x}}\left[\|\hat{\mathbf{x}}-\mathbf{u}\|^{2}_{2}\right]$ (17)
@@
Since $P(\mathbf{x})\geq 0$, minimizing the MSE is equivalent to minimizing the inner integral for each $\mathbf{x}$. Expanding the square and simplifying, we get:
[…]
Taking the derivative with respect to $\hat{\mathbf{x}}$ and setting it to zero, we find:
[…]
Thus, the posterior mean minimizes the MSE for any $\mathbf{x}$, and therefore minimizes the overall MSE.
Tweedie’s Formula and the MMSE Denoiser: While the MMSE expectation integral is generally difficult or impossible to evaluate directly, a key result known as Tweedie’s formula [31, 32, 33] enables us to write the expression for MMSE also in the form of the gradient of a scalar function:
@@
Data-driven MMSE denoisers: The typical modern supervised approach to image denoising is to train a deep neural network with pairs of clean and noisy images, where the noise is often modeled as additive white and Gaussian(AWGN) [10]. Let’s assume we have image pairs $(\mathbf{u},\mathbf{x})\sim P(\mathbf{u},\mathbf{x})$ where $\mathbf{u}$ represents a clean image, and $\mathbf{x}$ is the noisy observation obtained by adding AWGN with a known standard deviation to $\mathbf{u}$: $\mathbf{x}=\mathbf{u}+\mathbf{e}$, where $\mathbf{e}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$.
A typical regression approach would predict $\mathbf{u}$ directly from $\mathbf{x}$ using a trained model $\hat{\mathbf{x}}=F_{\theta}(\mathbf{x})\approx\mathbf{u}$, by minimizing the expected reconstruction error:
$\min_{\theta}\mathbb{E}_{\mathbf{u},\mathbf{e}}\|F_{\theta}(\mathbf{u}+\mathbf{e})-\mathbf{u}\|^{p}_{p}\approx\min_{\theta}\sum_{i}\|F_{\theta}(\mathbf{u}^{i}+\mathbf{e}^{i})-\mathbf{u}^{i}\|^{p}_{p}.$ (23)
@@
As mentioned earlier, the MMSE denoiser is the average of all plausible clean signals given the noisy observation. This averaging can lead to a loss of details and a blurry appearance, especially when the noise level is high. This is because minimizing average distortion (e.g., PSNR) can harm perceptual quality [34]. To address this, alternatives including perceptual [35, 36, 37] and adversarial losses [38, 39, 40] have been considered. A more powerful approach is to sample from the posterior distribution, avoiding the regression to the mean effect  [41, 42, 43, 44, 45, 46, 47, 48, 49].
@@ 3.2 Energy-based Denoisers
We’ve seen that both MMSE and MAP estimators are of the form $f(\mathbf{x},\alpha)=\mathbf{x}-\alpha\nabla\phi_{\alpha}(\mathbf{x})$ where $\phi_{\alpha}(\mathbf{x})$ is some smoothed version of $\phi$ – differently smoothed in each case. These denoisers are also special cases of a general “energy-based” formulation [51]:
$f(\mathbf{x},\alpha)=\nabla\mathcal{E}(\mathbf{x},\alpha),$ (24)
@@
$\mathcal{E}(\mathbf{x},\alpha)=\frac{1}{2}\|\mathbf{x}\|^{2}-\alpha\phi_{\alpha}(\mathbf{x}).$ (25)
@@
$\nabla f(\mathbf{x},\alpha)=\mathcal{H}\left[\mathcal{E}(\mathbf{x},\alpha)\right],$ (26)
@@
Approximation of Energy-based Denoisers
The energy-based formulation of denoisers provides a natural mechanism for principled empirical design of denoisers. This approach turns out to be consistent with the well-established empirical Bayes [31, 52] approach as well.
@@
$s(\mathbf{x},\alpha)=\nabla\log P(\mathbf{x},\alpha)=-\nabla\mathcal{E}(\mathbf{x},\alpha).$ (28)
@@
$s(\mathbf{x},\alpha)=-\nabla\mathcal{E}(\mathbf{x},\alpha)=-\nabla\mathcal{E}(0,\alpha)-\mathbf{L}(\mathbf{a},\alpha)\mathbf{x}.$ (30)
Meanwhile, Tweedie’s formula implies that the MMSE denoiser has the form:
$f(\mathbf{x},\alpha)$ $=\mathbf{x}+\alpha\>s(\mathbf{x},\alpha)$ (31)
$=\mathbf{x}-\alpha\nabla\mathcal{E}(0,\alpha)-\alpha\mathbf{L}(\mathbf{a},\alpha)\mathbf{x}.$
@@
$f(\mathbf{x},\alpha)=\mathbf{W}(\mathbf{x},\alpha)\>\mathbf{x},$ (33)
@@
Though these facts were neither historically clarified, nor the original motivation for their development [53], denoisers of the form (33) have always been heuristic/empirical approximations to the MMSE. These denoisers were hugely popular and effective (e.g. [56, 14]) for decades before the more recent introduction of neural networks. More recent work by Scarvelis et al.  [57] explores the use of a specific kernel approach to create a “closed-form” diffusion model that operates directly on the training set, without the need for training a neural network.